# PeRFlow: Piecewise Rectified Flow as Universal Plug-and-Play Accelerator

### v1

@@ 3.3 Universal acceleration to SD workflows
PeRFlow-$\Delta W$ serves as a universal accelerator that can be simply plugged into various pipelines based on the pretrained Stable Diffusion models including ControlNet [Zhang and Agrawala, 2023], IP-Adaptor [Ye et al., 2023], and multiview-generation. For example, we plug PeRFlow-$\Delta W$ into the SD-v1.5 ControlNet-Tile, this enables a 4-step image enhancement module, see figure 10. Using together with the 4-step PeRFlow T2I, we can generate high-quality $1024\times 1024$ images with lightweight SD-v1.5 backbones. For multiview-generation, we plug PeRFlow-$\Delta W$ into the Wonder3D [Long et al., 2023] pipeline. Surprisingly, we can generate clear multiview images with only one step, see figure 11. More results are shown in figure 13, 12.
3.4 Inference budget allocation and dynamic CFG
@@
Suppose PeRFlow divides the entire sampling time range into $K$ windows, $K$-step inference (one for each window) will yield high-quality images in most cases. For some pictures with complex structures, like motorcycles with well-crafted wheels and engines, PeRFlow may require more steps. Ho et al. [2020] found that diffusion models generate images by synthesizing the layout and structure first and then refining the local details. Inspired by this observation, we allocate the extra required steps to windows in highly noisy regions. Specifically, we have the $K$ time windows $\{[t_{k},t_{k-1})\}_{k=K}^{1}$ with $01=t_{K}>\dots>t_{k}>t_{k-1}>\dots>t_{0}=$ 0. Let the number of inference steps be $N<=2K$, we allocate $N//K+1$ steps for window $[t_{k},t_{k-1})$ if $K-k<N\%K$, otherwise $N//K$. In practice, PeRFlow creates $4$ time windows for acceleration training, and $5$-step inference can consistently generate high-quality images.
@@
CFG is a useful technique to improve the layout, structure, and text alignment of the generated images. However, a large CFG scale sometimes leads to over-saturated color blocks [Kynkäänniemi et al., 2024, Wang et al., 2024b]. To mitigate this issue, we use a dynamic CFG strategy for few-step sampling, i.e., the corresponding CFG scales decrease for window $K$ to $1$. For example, when sampling with $5$ steps, the CFG schedule is $7.5$-$4.0$-$4.0$-$4.0$ for the CFG-sync mode and $2.5$-$1.5$-$1.5$-$1.5$ for the CFG-fixed mode.
@@
Figure 10: Fast (4-step) image refinement ($128\times 128$ to $1024\times 1024$) via PeRFlow-ControlNet-tile.
Figure 11: PeRFlow works compatibility with Wonder3D.
@@
Figure 12: Fast generation via PeRFlow accelerated depth-/edge-/pose-ControlNet
Figure 13: PeRFlow works compatibility with IP-Adaptor.

### v3

@@ 3.2 PeRFlow as Universal Plug-and-Play Accelerator on SD Work Flows
PeRFlow-$\Delta W$ serves as a universal accelerator that can be simply plugged into various pipelines trained on the pretrained Stable Diffusion models, including (but not limited to) ControlNet [Zhang and Agrawala, 2023], IP-Adaptor [Ye et al., 2023], and multiview generation. For example, plugging PeRFlow-$\Delta W$ into the SD-v1.5 ControlNet-Tile gives a 4-step image enhancement module (figure 7). Combining this module with the 4-step PeRFlow-SD-v1.5, we can generate high-quality $1024\times 1024$ images with lightweight SD-v1.5 backbones. For multiview generation, plugging PeRFlow-$\Delta W$ into the Wonder3D [Long et al., 2023] pipeline leads to one-step generation of multi-view images (figure 8). More results are shown in figure 9.
3.3 Additional Discussion
@@
Inference Budget Allocation   If PeRFlow divides the entire sampling trajectory into $K$ windows and perfectly straightens the sub-trajectories in each window, $K$-step inference (one for each window) will yield high-quality images. However, for pictures with complex structures, such as motorcycles with well-crafted wheels and engines, PeRFlow may require more steps. Ho et al. [Ho et al., 2020] found that diffusion models generate images by synthesizing the layout and structure first and then refining the local details. Inspired by this observation, we allocate the extra steps to windows in highly noisy regions. Supposing we have $K$ time windows $\{[t_{k},t_{k-1})\}_{k=K}^{1}$ with $01=t_{K}>\dots>t_{k}>t_{k-1}>\dots>t_{0}=$ 0, we can allocate more than one steps to the earlier windows $\{[t_{k},t_{k-1})\}_{k=K}^{K_{extra}}$, where $K_{extra}$ is pre-defined. In practice, PeRFlow creates $4$ time windows for acceleration training, and $5$-step inference (with one extra step in $[t_{K},t_{K-1})$) consistently generates high-quality images.
@@
Dynamic Classifier-Free Guidance   CFG is a useful technique to improve the layout, structure, and text alignment of the generated images. However, a large CFG scale sometimes leads to over-saturated color blocks [Kynkäänniemi et al., 2024, Wang et al., 2024b]. To mitigate this issue, we use a dynamic CFG strategy for few-step sampling, i.e., the corresponding CFG scales decrease for window $K$ to $1$. For example, when sampling with $5$ steps, the CFG schedule is $7.5$-$4.0$-$4.0$-$4.0$ for the CFG-sync mode and $2.5$-$1.5$-$1.5$-$1.5$ for the CFG-fixed mode.
4 Related Works
Few-Step Diffusion Models   Diffusion models have demonstrated impressive generative capabilities, but their iterative sampling process often suffers from slow inference speed [Ho et al., 2020, Song et al., 2021, 2022]. To accelerate these models, various methods have been proposed. Progressive Distillation [Salimans and Ho, 2022, Meng et al., 2023] iteratively reduces the number of inference steps to 4-8, but the error can accumulate during the process. Alternative approaches [Xiao et al., 2021, Wang et al., 2022, Zheng et al., 2022, Xu et al., 2023a, Lin et al., 2024, Sauer et al., 2023] leverage adversarial losses to align the distributions and reduce the number of inference steps, but these methods often struggle with training instability and mode collapse. To avoid adversarial training, recent works [Yin et al., 2023, Nguyen and Tran, 2024, Zhou et al., 2024] employ additional models to estimate the score of the generated data for distilling one-step generators, but this adds extra cost to the training pipeline. Consistency Distillation [Song et al., 2023, Luo et al., 2023a] is a novel pipeline for distilling few-step diffusion models by optimizing a consistency loss. However, the substantial difference between consistency models and the original diffusion models can hurt their adaptability to pre-trained modules. In our work, PeRFlow provides a simple, clean, and efficient framework for training few-step generative flows. By using different parameterizations as described in Section 2.2, PeRFlow achieves minimal gap with diffusion models, making it suitable for various pre-trained workflows.
Straight Probability Flows   Learning straight probability flow is a promising principle for obtaining fast generative flows [Liu et al., 2022, Liu, 2022, Liu et al., 2023, Finlay et al., 2020]. Reflow is an effective way to learn such straight flows, but it requires constructing a large synthetic dataset [Liu et al., 2022, 2023], which can introduce computational overhead and distribution shift. To avoid dataset construction,[Lee et al., 2023, Xing et al., 2023] use an extra neural network to estimate the initial noise corresponding to an image, but training this network can be challenging. [Pooladian et al., 2023] employs mini-batch optimal transport to directly learn a straighter trajectory, but it is unclear how to apply this method to conditional generation scenarios, such as text-to-image generation. [Nguyen et al., 2023] finds the best step-size schedule for the pretrained generative model before reflow to improve efficiency, but it cannot avoid dataset generation and the resulting distribution shift. PeRFlow provides a new method to avoid using synthetic datasets. It uses real training data to mitigate distribution shift and a divide-and-conquer strategy to efficiently perform reflow, leading to advanced few-step text-to-image generators. Sequential reflow[Yoon and Lee, 2024] is a concurrent work to ours. Compared to their work, we additionally provide different parameterization strategies to enhance the empirical performance in accelerating pre-trained text-to-image models.
5 Conclusions
In this work, we present Piecewise Rectified Flow (PeRFlow), a novel technique to learn few-step flow-based generative models. PeRFlow adopts a divide-and-conquer strategy, separating the generation trajectory into intervals and applying the reflow operation within each interval. This yields two key advantages: (1) using real training data to mitigate distribution shift from synthetic data, and (2) avoiding the need to generate and store a synthetic dataset prior to training. PeRFlow also designs proper parameterizations to inherit knowledge from pre-trained diffusion models for fast convergence. Consequently, PeRFlow accelerates powerful diffusion models like SD v1.5, SD v2.1, and SDXL, producing high-quality few-step image generators. Moreover, PeRFlow can be seamlessly combined with various SD workflows to create their accelerated versions.
Limitations   Currently, PeRFlow divides the time range into 4 windows, balancing inference and training costs. It needs 4 steps or more for generation. To enable 1-2 step inference, we plan to explore multi-stage training and will focus on avoiding target synthesizing in the future.
Appendix
Appendix A More generation results
@@
Figure 7: 4-step image enhancement ($128\rightarrow 1024$) with PeRFlow-SD v1.5+ControlNet-tile [Zhang and Agrawala, 2023]
Figure 8: One-step multiview generation of PeRFlow-SD v1.5+Wonder3D [Long et al., 2023]
@@
Figure 9: Fast generation via PeRFlow accelerated depth-/edge-/pose-ControlNet [Zhang and Agrawala, 2023]
Figure 10: 4-step generation ($512\times 512$) via PeRFlow-SD-v1.5.
Figure 11: 8-step generation ($512\times 512$) via PeRFlow-SD-v1.5.
Figure 12: 4-step generation ($768\times 768$) via PeRFlow-SD-v2.1.
Figure 13: 4-step generation ($1024\times 1024$) via PeRFlow-SDXL.